Use chain rule to find the derivative and express the final answer in terms of x in radical form of

y= u1/2 and u= x1/2

1
Expert's answer
2021-09-20T16:02:23-0400

Let us use the chain rule "y'_x=y'_uu'_x:"

"y'_x=\frac{1}2u^{-\frac{1}2}\frac{1}2x^{-\frac{1}2}=
\frac{1}4(x^{\frac{1}2})^{-\frac{1}2}x^{-\frac{1}2}"

"=\frac{1}4x^{-\frac{1}4}x^{-\frac{1}2}
=\frac{1}4x^{-\frac{3}4}=\frac{1}{4\sqrt[4]{x^3}}."